# MotionBooth: Motion-Aware Customized Text-to-Video Generation

### v1

@@ 3.2 Subject Learning
Given a few images of a subject, previous works have demonstrated that fine-tuning a diffusion model on these images can effectively learn the appearance of the subject [39, 23, 8, 10, 40, 44]. However, two significant challenges remain. First, due to the limited size of the dataset, the model quickly overfits the input images, including their backgrounds, within a few steps. This overfitting of the background impedes the generation of videos with diverse scenes, a problem also noted in previous works [39, 12]. Second, fine-tuning T2V models using images can impair the model’s inherent ability to generate videos, leading to severe background degradation in the generated videos. To illustrate these issues, we conducted a toy experiment. As depicted in Fig. 3, without any modifications, the model overfits the background to the subject image. To address this, we propose computing the diffusion reconstruction loss solely within the subject region. However, even with this adjustment, the background in the generated videos remains over-smoothed. This degradation likely results from tuning a T2V model exclusively with images, which damages the model’s original weights for video generation. To mitigate this, we propose incorporating video data as preservation data during the training process. Although training with video data but without subject region loss still suffers from overfitting, our approach, MotionBooth, can generate videos with detailed and diverse backgrounds. (xiangtai: hard to understand.)
@@ 4.1 Experimental Setup
Evaluation metrics. We evaluate motion-aware customized video generation from three aspects: region subject fidelity, temporal consistency, and camera motion fidelity. 1) To ensure the subject is well-preserved and accurately generated in the specified motion, we introduce region CLIP similarity (R-CLIP) and region DINO similarity metrics (R-DINO). These metrics utilize the CLIP [37] and DINOv2 [35] models to compute the similarities between the subject images and frame regions indicated by bounding boxes. Additionally, we use CLIP image-text similarity (CLIP-T) to measure the similarity between entire frames and text prompts. 2) We evaluate temporal consistency by computing CLIP image features between each consecutive frame. 3) We use VideoFlow [42] to predict the optical flow of the generated videos. Then, we calculate the flow error by comparing the predicted flow with the ground-truth camera motion provided in the evaluation dataset.
@@
Method Module Weight Storage CLIP-T $\uparrow$ T-Cons. $\uparrow$ Flow error $\downarrow$
AnimateDiff [13] 74M 0.245 0.925 1.683
CameraCtrl [17] 2.5G 0.237 0.939 0.807
MotionBooth (Zeroscope) [No Training] 0.252 0.948 0.190
MotionBooth (LaVie) [No Training] 0.241 0.963 0.296
@@ Appendix A Appendix
A.1 Human Preference Study
A.2 Limitations and Future Work
Figure 9: Failure cases of MotionBooth.
A.3 Implementation Details

### v3

@@ 3.2 Subject Learning
Given a few images of a subject, previous works have demonstrated that fine-tuning a diffusion model on these images can effectively learn the appearance of the subject [42, 23, 8, 10, 43, 47]. However, two significant challenges remain. First, due to the limited size of the dataset, the model quickly overfits the input images, including their backgrounds, within a few steps. This overfitting of the background impedes the generation of videos with diverse scenes, a problem also noted in previous works [42, 12]. Second, fine-tuning T2V models using images can impair the model’s inherent ability to generate videos, leading to severe background degradation in the generated videos. To illustrate these issues, we conducted a toy experiment. As depicted in Fig. 3, without any modifications, the model overfits the background to the subject image. To address this, we propose computing the diffusion reconstruction loss solely within the subject region. However, even with this adjustment, the background in the generated videos remains over-smoothed. This degradation likely results from tuning a T2V model exclusively with images, which damages the model’s original weights for video generation. To mitigate this, we propose incorporating video data as preservation data during the training process. Although training with video data but without subject region loss still suffers from overfitting, our approach, MotionBooth, can generate videos with detailed and diverse backgrounds.
@@ 3.3 Subject Motion Control
Discussion. An important aspect is how the necessary information from other language tokens is integrated into the generated outputs, given that tokens such as verbs and background nouns are assigned minimal values. We propose that this information is extracted through the <start> and <end> tokens. Given that Transformer-based language encoders like CLIP [40] are typically trained on classification tasks, they often encode the overall context of a sentence into these special tokens. Thus, despite the suppression of other tokens during the softmax calculation, the model can still access relevant information about verbs, background elements, and other components necessary for the generation process. To support this explanation, we conducted an experiment in which we examined the softmax outputs using a naive text-to-video pipeline. The results showed that the <start> token consistently held the highest softmax value, close to 1, while the <end> token had the second-largest value. The remaining tokens, including those representing nouns, adjectives, verbs, and conjunctions, were distributed among the remaining softmax values.
@@ 4.1 Experimental Setup
Evaluation metrics. We evaluate motion-aware customized video generation from four aspects: region subject fidelity, temporal consistency, camera motion fidelity, and video quality. 1) To ensure the subject is well-preserved and accurately generated in the specified motion, we introduce region CLIP similarity (R-CLIP) and region DINO similarity metrics (R-DINO). These metrics utilize the CLIP [40] and DINOv2 [38] models to compute the similarities between the subject images and frame regions indicated by bounding boxes. Additionally, we use CLIP image-text similarity (CLIP-T) to measure the similarity between entire frames and text prompts. 2) We evaluate temporal consistency by computing CLIP image features between each consecutive frame. 3) We use VideoFlow [45] to predict the optical flow of the generated videos. Then, we calculate the flow error by comparing the predicted flow with the ground-truth camera motion provided in the evaluation dataset. 4) We randomly select 1000 videos from the MSRVTT dataset [64], predict their camera motion sequences with VideoFlow [45], and compute the FVD metric for camera motion control only.
@@
Method Module Weight Storage FVD $\downarrow$ CLIP-T $\uparrow$ T-Cons. $\uparrow$ Flow error $\downarrow$
Text2Video-Zero (SD 1.5) [26] [No Training] 1821.72 0.248 0.904 1.854
AnimateDiff [13] 74M 1515.82 0.245 0.925 1.683
CameraCtrl [17] 2.5G 1468.53 0.237 0.939 0.807
MotionCtrl [56] 4.0G 1109.45 0.236 0.935 0.872
MotionBooth (Zeroscope) [No Training] 905.40 0.252 0.948 0.190
MotionBooth (LaVie) [No Training] 723.26 0.241 0.963 0.296
@@ 4.3 Ablation Studies
4.4 Human Preference Study
4.5 Limitations and Future Work
Figure 9: Failure cases of MotionBooth.
@@ Appendix A Appendix
A.1 Implementation Details
@@
(a) Examples of enlarging or shrinking the subject bounding box.
(b) Examples of gradually enlarging the camera motion speed.
(c) Examples of camera motion control with large foreground objects.
(d) Example of combining zoom-in and camera motion control.
(e) Example of masking images during training.
(f) Ablations of the filling method in latent shift.
(g) Examples of controlling both subject and camera motion.
(h) Comparison of latent shift and text guidance to control camera motion.
Figure 12: More qualitative results.
A.2 Discussions
Differences with related works. Our proposed subject and camera motion control methods differ in several ways from previously established approaches, such as TrailBlazer [34], Directed Diffusion [35], Boximator [52], Motion-Zero [4], MotionCtrl [56], and Text2Video-Zero [26].
TrailBlazer [34] and Directed Diffusion [35] use a training-free approach for controlling object motion by manipulating the cross-attention maps. Specifically, TrailBlazer [34] adjusts both spatial and temporal cross-attention by scaling the attention maps with a hyper-parameter of less than 1. Directed Diffusion [35] focuses on image generation in a similar manner. Our approach, in contrast, targets only the spatial cross-attention, setting attention values outside the object’s bounding box to zero. This targeted manipulation simplifies implementation and significantly enhances the performance in generating motion-aware customized videos.
Boximator [52] relies on a training-based technique, requiring box coordinates as inputs to a newly trained self-attention layer. In comparison, our method is training-free, thus providing a more accessible and user-friendly solution for controlling subject and camera motion.
Motion-Zero [4] also operates without additional training, but it utilizes a test-time tuning technique to adjust the latent space using cross-attention map loss during the denoising process. This technique, however, increases the generation time and memory requirements considerably—from approximately 15 seconds to several minutes per video. Additionally, our experiments indicate that Motion-Zero often produces poor-quality outputs, with videos collapsing into unrecognizable elements. This outcome is likely due to the detrimental effects of test-time tuning on parameters and latent distributions in customized scenarios. In contrast, our method directly manipulates the cross-attention map, adding only 0.3 seconds to the generation process and consistently yielding more reliable visual outcomes.
MotionCtrl [56] uses a training-based approach that requires inputting point trajectories to control camera poses and object motion. Our method does not involve additional training, offering a simpler alternative for subject and camera motion control.
Text2Video-Zero [26] builds on a pre-trained text-to-image (T2I) model and extends it to video generation by utilizing consistent noise across frames. However, this approach is unsuitable for text-to-video (T2V) models, which typically use different noise for each frame. Additionally, Text2Video-Zero uses latent shifting for overall scene movement and mirrors latent information to fill in missing regions. In comparison, our approach employs random sampling along the x and y axes, resulting in more concise and coherent video generation for camera motion control.
We conducted quantitative experiments to evaluate the performance of these methods, as reported in Table 2 and Table 4(b). Some methods are not included due to the unavailability of their code. Nonetheless, our results indicate that our method generally outperforms the other alternatives. In particular, we observed significant limitations in the test-time tuning strategy of Motion-Zero [4] when applied to customized video generation, which further emphasizes the strengths of our approach.
Zoom-in/out effect with subject motion control. Our subject motion control technique effectively manages changes in bounding box size, such as gradual enlargement or reduction. Specifically, enlarging the bounding box results in an appropriate scaling of the subject, creating a zoom-in effect in the generated video. Conversely, reducing the bounding box size produces a zoom-out effect. This zoom-in and zoom-out behavior is demonstrated in Fig. 12(a), where examples illustrate these effects in response to bounding box adjustments. LABEL:{fig:zoom-in-camera-move} shows an example of combining zoom-in and camera motion control, demonstrating the method’s flexibility.
Large camera motion speeds. We evaluate the performance of our camera motion control technique under varying levels of camera movement intensity. As depicted in Fig. 12(b), the method was first tested with a camera movement vector of $[c_{x},c_{y}]=[-0.5,0.45]$, corresponding to a movement of half the video width to the left and nearly half downward. Under these conditions, our method successfully managed the camera movement, producing accurate results. When the camera motion speed was doubled, the model continued to perform well, demonstrating its ability to handle high-speed scenarios. However, at a speed three times the initial setting, i.e., $[c_{x},c_{y}]=[-1.5,1.35]$, the method exhibited limitations, resulting in only a downward tiling effect. These findings indicate that while our method can effectively manage camera movements spanning the entire video width, its performance diminishes under extremely high-speed conditions.
Camera motion control with large foreground objects. We evaluate the camera control capability in scenes containing substantial foreground objects. The results, presented in Fig. 12(c), show an experiment involving a scene dominated by a large candy house, which nearly occupies the entire frame. Despite the significant presence of the foreground object, our technique effectively managed to pan the camera to the right, suggesting that the method is capable of handling complex scenes with large foreground elements. This robustness is attributed to the method’s reliance on latent shifts, which simultaneously move both foreground and background elements, thereby ensuring that the presence of large foreground objects does not significantly impair performance.
Model efficiency. Training our model by fine-tuning a specific subject takes approximately 10 minutes. For LaVie [54], the naive text-to-video (T2V) pipeline takes about 15.0 seconds per video. When incorporating subject control, the inference time is 15.3 seconds per video; with camera control, it increases to 20.6 seconds per video; and when applying both camera and subject control, the inference time is 21.5 seconds per video.
A.3 Comparison with More Baselines
Table 4: Comparison with More Baselines.
(a) Comparison of the latent shift method and text guidance for camera motion control.
Method CLIP-T $\uparrow$ T-Cons. $\uparrow$ Flow error $\downarrow$
Text Guidance 0.256 0.957 0.416
Latent Shift (Ours) 0.252 0.948 0.190
(b) Comparison of subject motion control with more baselines.
Method R-CLIP R-DINO CLIP-T $\uparrow$ T-Cons. $\uparrow$
Motion-Zero [4] - - Collapse -
Directed Diffusion [35] 0.668 0.242 0.260 0.954
TrailBlazer [34] 0.669 0.251 0.259 0.957
Ours (Zeroscope) 0.767 0.305 0.242 0.968
Ours (LaVie) 0.735 0.247 0.244 0.965
Comparison with text guidance for camera motion control. we compare our latent shift method with text guidance for controlling camera motion with the results presented in Table 4(a) and  Fig. 12(h). Specifically, we utilize simple text prompts, such as “camera pan right” and “camera pan up right,” to influence camera movement. The generated videos reflect these instructions to some extent, especially for straightforward motions like panning to the right. However, we observe that such text prompts often lack sufficient specificity, particularly in terms of conveying critical details such as the speed or distance of the camera movement. As a result, this approach yielded sub-optimal outcomes when compared to our proposed method. This indicates that our approach offers a more precise and stable mechanism for controlling camera motion than is possible with text prompts alone.
Comparison of subject motion control with more baselines. Table 4(b) presents quantitative experiments comparing our method with several baseline approaches, including TrailBlazer [34], Directed Diffusion [35], and Motion-Zero [4]. The results demonstrate that our approach generally outperforms these alternatives.
@@ A.4 Ablation Studies
Table 6: More ablation studies.
(a) Ablation of controlling single motion type.
T2V Model Evaluation Set R-CLIP $\uparrow$ R-DINO $\uparrow$ CLIP-T $\uparrow$ T-Cons. $\uparrow$ Flow error $\downarrow$
Zeroscope Only Subject 0.767 0.305 0.242 0.968 -
Only Camera - - 0.252 0.948 0.190
Both 0.667 0.306 0.258 0.958 0.252
LaVie Only Subject 0.735 0.247 0.244 0.965 -
Only Camera - - 0.241 0.963 0.296
Both 0.712 0.472 0.247 0.962 0.332
(b) Ablation of masking the training images.
Method R-CLIP $\uparrow$ R-DINO $\uparrow$ CLIP-T $\uparrow$ T-Cons. $\uparrow$
Mask Image 0.683 0.060 0.237 0.937
Mask Loss (Ours) 0.712 0.472 0.247 0.962
(c) Ablation of the latent filling method in latent shift.
Method CLIP-T $\uparrow$ T-Cons. $\uparrow$ Flow error $\downarrow$
Random Init 0.235 0.922 0.212
Random Sample 0.248 0.954 0.185
Loop 0.252 0.956 0.168
Reflect - Collapse -
Ours 0.252 0.948 0.190
Controlling Single Motion Type. We conduct an additional experiment focusing specifically on the scenario where only subject motion is controlled, and the results are summarized in Table 6(a). Notably, we observe that the inclusion of both subject and camera motion controls leads to a slight decrease in some metrics. For instance, using the Zeroscope model, there is a 0.1 drop in the R-CLIP score and a 0.01 decrease in the T-Cons metric when both types of motion are controlled compared to the scenario where only subject motion is controlled. However, there are also instances where metrics such as R-DINO and CLIP-T show a slight improvement under combined control conditions. We consider this trade-off acceptable within the context of motion-aware customized video generation. Qualitative results in Fig. 12(g) show that our method can successfully handle controlling both subject and camera motion.
Masking the training images. We conduct an ablation study to evaluate the effect of masking the background regions directly in pixel space of the training images, compared to masking the diffusion loss. As illustrated in Fig. 12(e), directly masking the training images significantly impairs the model’s capacity to learn the subject effectively. The quantitative metrics shown in Table 6(b) indicate a marked decrease in performance when the training images are masked. We hypothesize that this decline arises from the unnatural distribution created by the masked images, which disrupts the learning process. Consequently, we conclude that masking the diffusion loss, rather than masking the training images directly, is a more effective strategy for preserving the integrity of the learned representations.
Latent Filling Methods. We test several latent filling approaches and present the qualitative results in Fig. 12(f) and the quantitative results in Table 6(c).
Random Init: This method involves filling the hole with random values. Our experiments revealed that this technique leads to severe artifacts due to the disruption of the natural horizontal and vertical distribution of latent pixels, ultimately degrading the overall visual quality.
Random Sample: In this approach, values are randomly sampled in the latents. Similar to Random Init, this method produced significant artifacts, leading to poor visual quality in the generated video.
Loop: This method reuses the moved-out-of-scene values to fill the missing region, thereby creating a looping background effect. While this technique was found to preserve video quality better than the random methods, it introduced a limitation in terms of flexibility for camera movements, resulting in repetitive looping effects. Therefore, it is not suitable for more diverse camera controls.
Reflect: This approach is employed in Text2Video-Zero [26], where the missing region is filled by reflecting the surrounding content. However, in our Text-to-Video (T2V) scenario, this method collapsed, failing to maintain the desired visual quality.
The quantitative results presented in Table 6(c) corroborate these findings. Random Init and Random Sample lead to significant artifacts, whereas the Loop method provides better visual quality but at the cost of limiting camera movement diversity. The Reflect method, despite its success in other applications, did not yield satisfactory results in our T2V context.
In conclusion, our proposed straight-sample method consistently maintained high visual quality without introducing significant artifacts or limiting camera movement flexibility, demonstrating its superiority in maintaining the desired video fidelity.